IR SPECTROSCOPY OF Au^{−}·(CO_{2})_{n} CLUSTERS: STRONG CLUSTER SIZE DEPENDENCE OF METAL-LIGAND INTERACTION

\underline{BENJAMIN J. KNURR} and J. MATHIAS WEBER, JILA, NIST and Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309.

Gold is a widely used catalyst in many reactions. For example, negatively charged gold clusters have been shown to catalyze the oxidation of CO to CO_{2} [1], although the precise role of negative charge has not been understood. A previous study from our laboratory has shown that the binary complex [AuCO_{2}]^{−} has the CO_{2} ligand covalently bound to the gold, leading to significant charge transfer onto the CO_{2} unit and concomitant decrease of the OCO bond angle and weakening of the CO bonds [2]. The structure of this aurylformate anion is reminiscent of structural motifs assumed to play a role in a recent approach towards production of solar fuels using reductive activation of CO_{2} [3].

We report infrared spectra of Au^{−}·(CO_{2})_{n} clusters highlighting solvation mediated changes in the infrared signatures of the ligands. These results are discussed in the framework of quantum chemistry calculations.

1. U. Heiz and W.D. Schneider, J. Phys. D - Appl. Phys., 33 (2000) R85- R102.

2. A.D. Boese, H. Schneider, A.N. Gloess, and J.M. Weber, J. Chem. Phys., 122 (2005) 154301.

3. E.E. Barton, D.M. Rampulla, and A.B. Bocarsly, J. Am. Chem. Soc., 130 (2008) 6342.